if a tube have filed with water to a height 5m.and suddenly air of specific density of 0.85started to flow over it with 50m/s.find the the new height of water in tube.
1
Expert's answer
2020-09-14T10:23:49-0400

By the law of energy conservation:

"mgh=\frac{mv^2}{2}+mgh^{'}"

"gh=\frac{v^2}{2}+gh^{'}"

"g(h-h{'})=\frac{v^2}{2}"

"h-h^{'}=\frac{v^2}{2g}"

"h^{'}=h-\frac{v^2}{2g}"

"h{'}=5-\frac{2500}{20}=-120"